OpenStudy (anonymous):

what is the slope of a line that passes through the point (-1, 1) and is parallel to a line that passes through (4, 6) and (-1,-4)?

OpenStudy (anonymous):

The slope of parallel lines is the same. So first find the equation of the line for (4,6) and (-1,-4) (-10)/(-5)=2 So we plug in for the point (-1,1): y-1=2(x+1) =2x+3

OpenStudy (anonymous):

so rise over run..wait is that the slope?

OpenStudy (anonymous):

Oh sorry, ya I didnt label my slope. slope is \[m=\frac{ y-y1 }{ x-x1 }\]

OpenStudy (anonymous):

So that was what I was doing with (-10)/(-5), plug in the points.

OpenStudy (anonymous):

right, so slope is -2/-1?

OpenStudy (anonymous):

ya so its 2 it doesnt matter what order you plug in the points

OpenStudy (anonymous):

Did it work?

OpenStudy (anonymous):

the slope is 2 or -2? I think it's working!

OpenStudy (anonymous):

@natalie135

OpenStudy (anonymous):

The slope is 2 since (-2)/(-1) simplifies to just 2.

OpenStudy (anonymous):

thank you so much for explaining! :)